OpenStudy (anonymous):

squareroot-15 squareroot-4 simplify using i

OpenStudy (goldeverything):

what is the question?

OpenStudy (anonymous):

\[\sqrt{-15} \sqrt{-4}\] simplify using i

OpenStudy (goldeverything):

ohhh ok

OpenStudy (anonymous):

root (15) i * 2i, as i is root(-1)

OpenStudy (anonymous):

I'm lost.. I don't understand i

OpenStudy (anonymous):

"i" is used in complex numbers to stand for "-1" as normally you can square root a negative number you use it to make it "positive" in a sense so \[\sqrt{-15}=\sqrt{15i ^{2}}\]

OpenStudy (anonymous):

so \[\sqrt{15i} \sqrt{4i}?\]

OpenStudy (anonymous):

@GoldEverything help.. lol

OpenStudy (goldeverything):

i will try to help

OpenStudy (anonymous):

Thank you(:

OpenStudy (goldeverything):

one second a bad song came on i have to change it

OpenStudy (anonymous):

lol okay

OpenStudy (anonymous):

now i'm lost

OpenStudy (anonymous):

sorry i meant "i" is \[\sqrt{-1}\]

OpenStudy (anonymous):

i^2 = -1 so when you square root a negative number, just tack on an i with it for example \[\sqrt{-4} = 2i\]

OpenStudy (anonymous):

sorry for all those mistakes on my side

OpenStudy (anonymous):

so what about the -15?

OpenStudy (goldeverything):

ok so basically its 15 square root and the you sqare 15 beacuse it was negative

OpenStudy (anonymous):

soooo \[\sqrt{15}\sqrt{4}?\]

OpenStudy (anonymous):

you can't factor the square root of 15, so just bring out an i\[\sqrt{-15} = i \sqrt{15}\]

OpenStudy (anonymous):

Then how do I put this together??

OpenStudy (goldeverything):

im not that great with this equation

OpenStudy (anonymous):

It's like chinese to me. Lol

OpenStudy (anonymous):

\[\sqrt{-4} = 2i \] \[\sqrt{-15} = i \sqrt{15}\] \[i \sqrt{15}*2i = 2i^{2}\sqrt{15}\] \[i^{2} = -1\] \[2i^{2}\sqrt{15}=2(-1)\sqrt{15} = -2\sqrt{15}\]

OpenStudy (anonymous):

How do you know all this robz8?

OpenStudy (anonymous):

practice :)

OpenStudy (goldeverything):

funny thing is a can speak chinese @Saralynnjane

OpenStudy (anonymous):

Youre my new math person. Lol. And @GoldEverything That's awesome! I can speak a little polish.

OpenStudy (goldeverything):

i can speak some too not much though

OpenStudy (anonymous):

Where are you from then?

OpenStudy (goldeverything):

i was born in us but my mom and i have moved around alot in asia then we came back

OpenStudy (anonymous):

Ohhh. that explains a lot! Ive only lived in the usa.

OpenStudy (goldeverything):

yeah i learned most of those for school

OpenStudy (anonymous):

Military bases?

OpenStudy (goldeverything):

no

OpenStudy (anonymous):

hmmm okay

OpenStudy (goldeverything):

just normal places in asia

OpenStudy (goldeverything):

how much polish do you speak

OpenStudy (anonymous):

not very well lol